OpenStudy (anonymous):

Verify the trigonometric identity. cot(x)sec^4(x)=cot(x)+2tan(x)+tan^3(x)

OpenStudy (anonymous):

\[\cot x\sec^4x=\cdots\\ \cot x\left(\sec^2x\right)^2=\cdots\\ \cot x\left(\tan^2x+1\right)^2=\cdots\\ \cot x\left(\tan^4x+2\tan^2x+1\right)=\cdots\] Distribute the \(\cot x\) and you're done.

OpenStudy (anonymous):

LAW OF SINES Solve the triangle. B = 73°, b = 15, c = 10

OpenStudy (anonymous):

that's what it says lol